a circular swimming pool is surrounded by a walkway that is 4 ft wide. The circumference of the outer edge of the walkway is 190 ft. what is the radius of the pool to the nearest foot ?
Answer by mananth(16946)

Circumference of outer edge = 190 ft
2*pi*r=190
r= 190/2*pi
r= 95/*pi
r= 30.24
width = 4 ft
radius of pool = 30.24-4
=26 feet